neutrino scattering cross section σ ∼ 10^{−47}m^{2}

volume swept out V_{s} = Lσ mass of cube M = ρV = ρ (1m^{3})

number of nucleons per cubic meter N = \frac{M}{m_{p}} = \frac{ρ}{m_{p}}

volume per nucleon V_{N} = \frac{m_{p}}{ρ}

volume swept out V_{s} = volume per nucleon V_{N}
Lσ = \frac{m_{p}}{ρ}
L = \frac{m_{p}}{ρσ}